OpenStudy (anonymous):

Please help me with this... my head hurts Solve –--(-3x^2 – 4x – 4 = 0) ?????

OpenStudy (anonymous):

Do you know what the general formula for quadratic equations is?

OpenStudy (anonymous):

a^2 +b^2 =C^2

OpenStudy (tkhunny):

Please multiply by -1 before you do anything else.

OpenStudy (anonymous):

is it completely nescesary??

OpenStudy (anonymous):

ignore the three negatives before the equation

OpenStudy (anonymous):

No, it would be \[-b \pm \frac{ \sqrt{b^2-4ac} }{ 2a } \]

OpenStudy (anonymous):

hmmm okay, so I plug the values in?

OpenStudy (anonymous):

where a=-3, b=-4, and c=-4. So just plug in

OpenStudy (anonymous):

okay let me try, hold on please...

OpenStudy (tkhunny):

Not quite. A little format problem, there. \(x=\dfrac{-b \pm \sqrt{b^{2}- 4ac}}{2a}\)

OpenStudy (anonymous):

Exactly, what @tkhunny said. im sorry, im getting too old for computers and stuff

OpenStudy (anonymous):

is it 6 Root 2

OpenStudy (anonymous):

I didnt really do the problem,but i would expect you to get two sets of answers

OpenStudy (anonymous):

I think I got the right answer, if not oh well, thanks for the help man

OpenStudy (anonymous):

No problem. Glad to help